OpenStudy (anonymous):

The area of the cross-section of a prism is 45cm squared. The volume of the prism is 405cm cubed. Work out the length of the prism. Will medal if clear explanation is provided

OpenStudy (cggurumanjunath):

405/45

OpenStudy (anonymous):

@ParthKohli

OpenStudy (cggurumanjunath):

9CM

OpenStudy (anonymous):

can you please draw the shape out

OpenStudy (cggurumanjunath):

V=LBH =A*L ; LB=A

OpenStudy (cggurumanjunath):

V=A*L 405=45*L L=405/45=9CM

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

@SyedMohammed98

TheSmartOne (thesmartone):

For example, lets say that this prism is a rectangular prism, then the cross section's area would be the area of the traingle. And it will have the same cross section all along its length. 

OpenStudy (anonymous):

yep

TheSmartOne (thesmartone):

So the area of the cross section times the length = volume of the prism

TheSmartOne (thesmartone):

and they told us area of the cross section = 45 volume = 405

OpenStudy (anonymous):

please can you add the values on the diagram so its easier for me to visualise

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):